OpenStudy (anonymous):

Find equations of both the tangent lines to the ellipse x2 + 4y2 = 36 that pass through the point (12, 3).

myininaya (myininaya):

find y'

myininaya (myininaya):

2x+8yy'=0

myininaya (myininaya):

8yy'=-2x y'=-2x/(8y) y'=-x/(4y)

myininaya (myininaya):

now plug in (12,3) to figure out the slope of the tangent line

OpenStudy (jamesj):

I distinctly remember doing this earlier today . Obviously gogetta you're in the same class as someone else! Or maybe same text book!

myininaya (myininaya):

then use (12,3) and y' evaluated at (12,3) to find the equation of the tangent line

myininaya (myininaya):

y'=-12/(4 *3)=-1 is slope

OpenStudy (anonymous):

yeah i made a mistake!

myininaya (myininaya):

\[y=-1x+b \] \[3=-1(12)+b => b=15 \]

myininaya (myininaya):

y=-x+15

OpenStudy (anonymous):

@myininaya you made a mistake as well

OpenStudy (anonymous):

\[(12,3)\] is not on the graph. you have to find two lines tangent to the graph that pass through \[(12,3)\]

myininaya (myininaya):

omg that point is not even on the graph lol

myininaya (myininaya):

yes i just checked lol

OpenStudy (jamesj):

No -- you've solved assuming that (12,3) lies on the ellipse. But it most definitely does not, as that point doesn't satisfy the ellipse equation. So, the question is a little more subtle.

myininaya (myininaya):

i didn't read the question lol

OpenStudy (anonymous):

i think you need to reason that if the point (x,y) is on the graph then \[\frac{y-3}{x-12}=-1\] and also that \[x^2+4y^2=36\] and solve from there

OpenStudy (anonymous):

@ jamesj apparently you solved this already, so instead of me messing up the algebra maybe you can link to it

OpenStudy (jamesj):

Hmm ... it was about 8 hours ago. Let's just do it again. Note first that it is not necessarily true that slope of the line will be -1

OpenStudy (jamesj):

Draw a diagram and convince yourself of that.

OpenStudy (anonymous):

oh right the whole thing is wrong. in fact the slope cannot be -1

OpenStudy (jamesj):

What we need to do is write down the general equation for a tangent line and find those that pass through (12,3)

OpenStudy (anonymous):

should be \[\frac{y-3}{x-12}=-\frac{x}{4y}\]

OpenStudy (jamesj):

Right AND we know the x and y satisfy the ellipse equation. Now solve.

OpenStudy (anonymous):

i'll leave that as an exercise...

OpenStudy (jamesj):

lol

OpenStudy (anonymous):

i must have messed this up somewhere. think my algebra is gone

OpenStudy (jamesj):

Let's just do it. Multiplying out 4y^2 - 12y = -x^2 + 12x x^2 + 4y^2 = 12(x + y) Now LHS = 36 so x + y = 3

OpenStudy (anonymous):

yeah i got that far...

OpenStudy (anonymous):

ok now i see . we can write \[y=3-x\] and put \[x^2+4(3-x)^2=36\] so get y

OpenStudy (jamesj):

Something's wrong.

OpenStudy (phi):

While you guys figure out the hard tangent, the easy one is y= 3

OpenStudy (jamesj):

There's a mistake somewhereI can't put my finger on it.

OpenStudy (anonymous):

what i am wondering is when you solve \[\frac{y-3}{x-12}=-\frac{x}{4y}\] and \[x^2+4y^2=36\] don't you get an infinite number of solutions? it just comes out to \[x+y=3\]

OpenStudy (anonymous):

oh wait...

OpenStudy (anonymous):

\[x^2+4(3-x)^2=36\] \[5x^2-24x=0\] \[x=0,x=\frac{24}{5}\] whew

OpenStudy (jamesj):

Yes, but that's problematic, because the tangents at x = 0 are the lines y = 2, y =-2

OpenStudy (jamesj):

So what about those at x = 24/5 ... ah yes, those work.

OpenStudy (phi):